OpenStudy (anonymous):

Question #1) Find (a) the elasticity of demand and (b) the range of prices for which the demand is elastic (E , -1). f(p) = 200(30 - p) Question #2) Suppose that at price p = 15 dollars the demand for a product is elastic. If the price is raised, what will happen to revenue?

OpenStudy (anonymous):

@campbell_st can u help?

OpenStudy (campbell_st):

sorry... its not something I'm familar with... and I've just done a quick read on google... I'd need a fair bit of time to get my head around it. Sorry I can't help.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@dpaInc can u help?

OpenStudy (campbell_st):

here is a bit of information I just found... it may help http://www.mbs.edu/home/jgans/mecon/value/Popups/pop_up_examples_calculating_elasticity.htm

OpenStudy (campbell_st):

and here is another... http://www.extension.iastate.edu/agdm/wholefarm/pdf/c5-207.pdf

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

so reading the information if f(p) = 200(30 - p) and p = 10 f(10) = 4000 p = 20 f(20) = 2000 then the elasticity is \[\frac{\frac{4000 -2000}{4000+2000}}{\frac{10 - 20}{10+20}} = \frac{\frac{1}{3}}{\frac{-1}{3}}\] so it would appear to be - 1 thats a best guess...and 5 mins prep...

OpenStudy (anonymous):

i'll look at the information more

OpenStudy (anonymous):

ok so for question #2 i'll have to substitute 15 right?

OpenStudy (anonymous):

can u help with the second q? @Mertsj

OpenStudy (mertsj):

I would say you are right.

OpenStudy (anonymous):

ok thx